An ore sample contains 11% impurity in additon to a mix of CuS and Cu2S. Heating 100g of the mix produces 75.4g of copper metal with a purity of 89.5%. What is weight 5 of CuS and Cu2S in the ore?
1
Expert's answer
2012-09-14T09:53:32-0400
75.4* 0.895=67.483 - it is the mass of pure copper
mass of S in sample is
100-67.483-100*0.11=21.517
n of S= 21.517/32=0.6724 moles
n of Cu
=67.483/63.5=1.06 moles
64* (1.06 - x) + 32*(0.64-x) + 64*2*x+ 32*x =
21.517+67.483
67.84-64x+20.48-32x+128x+32x=89
64x=0.68

x=0.01

w of CuS=0.01/1.06*100=1%
m of CuS=89*0.01=0.89g
m of
CU2S=88.11 g